# Quasi-Bayes empirical Bayes: a sequential approach to the Poisson compound decision problem

### v1

@@ C Proofs for Section 2
We can write that
$\sum_{y\in\mathbb{N}_{0}}(\hat{\theta}_{g_{n}}(y)-\hat{\theta}_{g^{*}}(y))^{2}p_{g^{*}}(y)$
$=\sum_{y\in\mathbb{N}_{0}}\left(\frac{\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})}-\frac{\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}\right)^{2}\frac{1}{y!}\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})$
$\leq 2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\left(\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}\left(g_{n}(\vartheta_{i})-g^{*}(\vartheta_{i})\right)\right)^{2}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}$
$+2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\left(\sum_{i=1}^{d}\vartheta_{i}^{y+1}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}}{\left(\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g_{n}(\vartheta_{i})\right)^{2}\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}\left(\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}(g_{n}(\vartheta_{i})-g^{*}(\vartheta_{i})\right)^{2}$
$\leq 2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\sum_{i=1}^{d}\vartheta_{i}^{2y+2}e^{-2\vartheta_{i}}||g_{n}-g^{*}||^{2}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}+2\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\vartheta_{d}^{2}\sum_{i=1}^{d}\vartheta_{i}^{2y}e^{-2\vartheta_{i}}||g_{n}-g^{*}||^{2}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}$
$\leq 4\vartheta_{d}^{2}||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}\sum_{y\in\mathbb{N}_{0}}\frac{1}{y!}\frac{\sum_{i=1}^{d}\vartheta_{i}^{2y}e^{-2\vartheta_{i}}}{\sum_{i=1}^{d}\vartheta_{i}^{y}e^{-\vartheta_{i}}g^{*}(\vartheta_{i})}$
$\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2},$
with $C=4d\vartheta_{d}^{2}e^{\vartheta_{d}^{2}/\vartheta_{1}-2\vartheta_{1}+\vartheta_{d}}$. Thus,
${\rm Regret}(g_{n},g^{*})\leq C||\mathbf{g}_{n}-\mathbf{g}^{*}||^{2}.$
@@
Under the assumptions of Theorem 2.4, as $n\rightarrow\infty$ the stochastic approximation in $\Delta^{d}$:
@@
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)-\bf 1\right),$
${\bm{\epsilon}}_{n+1}=\mathbf{g}_{n}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}_{n}}(Y_{n+1})}-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}_{n}}(y)}p_{\mathbf{g}^{*}}(y)\right),$
and initial state $\mathbf{g}_{0}\in\mathring{\Delta}^{d}$, converges to $\mathbf{g}^{*}$, $\mathds{P}^{*}$-a.s.
@@
Let $J=\partial\Delta^{d}\cup\{g^{*}\}$. We first prove that ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$, $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$, where ${\rm d}$ denotes the Euclidean distance in $\mathbb{R}^{d-1}$. The result follows from Theorem 2.2.1′′ in Chen (2005), if we can prove that the following conditions hold:
  • A1
@@
  • A2
    $J$ is closed, and there exists a is continuously differentiable function $\mathscr{l}:\Delta^{d}\rightarrow\mathbb{R}$ such that $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},J)\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$:
  • A3
    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}\epsilon_{i+1}||=0,$ $\mathds{P}^{*}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$;
  • A4
    $\mathbf{h}$ is measurable and locally bounded.
Let $\mathscr{l}$ be defined as
$\mathscr{l}(\mathbf{g})=\sum_{y\in\mathbb{N}_{0}}\log\left(\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}\right)p_{\mathbf{g}^{*}}(y).$
A1 holds by assumption; A3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and A4 is obvious.
We now prove A2. The set $J$ is closed and $\mathscr{l}$ is continuously differentiable. Moreover, for $\mathbf{g}\in\mathring{(\Delta^{d})}$,
$\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}$ $=-\left(\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)-\bf 1\right)\right)^{T}\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}-\mathbf{k}_{y}(\vartheta_{d})\mathbf{1}}{p_{\mathbf{g}}(y)}p_{\mathbf{g}^{*}}(y)\right)$
$=-\left(\sum_{i=1}^{d}\mathbf{g}_{i}\left(\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}-1\right)^{2}\right)$
$\leq-\left(\sum_{i=1}^{d}\mathbf{g}_{i}\left(\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}-1\right)\right)^{2}=0,$
where, in the inequality, we have used Jensen inequality. We now prove that the inequality is strict, unless $\mathbf{g}=\mathbf{g}^{*}$. By contradiction, assume that there exist $\mathbf{g}^{\prime}$ and $\mathbf{g}^{\prime\prime}$ with $\mathbf{g}^{\prime}\neq\mathbf{g}^{\prime\prime}$ such that $\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}}(y)}=1$ for every $i=1,\dots,d$ both for $\mathbf{g}=\mathbf{g}^{\prime}$ and for $\mathbf{g}=\mathbf{g}^{\prime\prime}$. Then, for every $a\in(0,1)$
$1=\sum_{i=1}^{d}(a\mathbf{g}^{\prime}_{i})+(1-a)\mathbf{g}^{\prime\prime}_{i})\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{g^{*}}(y)}{ap_{\mathbf{g}^{\prime}}(y)+(1-a)p_{\mathbf{g}^{\prime\prime}}(y)}$
$<\sum_{i=1}^{d}(a\mathbf{g}^{\prime}_{i}+(1-a)\mathbf{g}^{\prime\prime}_{i})\left(a\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}^{\prime}}(y)}+(1-a)\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{\mathbf{g}^{*}}(y)}{p_{\mathbf{g}^{\prime\prime}}(y)}\right)$
@@
which leads to a contradiction. Hence A2 holds, and ${\rm d}(\mathbf{g}_{n},J)\rightarrow 0$ $\mathds{P}^{*}$-a.s., as $n\rightarrow\infty$.
To conclude, we need to show that $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{d})=0)=0$. Since ${\rm d}(\mathbf{g}^{*},\partial\Delta^{d})>0$ and ${\rm d}(\mathbf{g}_{n},\{\mathbf{g}^{*}\}\cup\partial\Delta^{d})\rightarrow 0$ $\mathds{P}-a.s.$, then $\mathds{P}^{*}(\liminf_{n}{\rm d}(\mathbf{g}_{n},\partial\Delta^{d})=0)>0$ if and only if there exists $M$ with $\mathds{P}^{*}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),\partial\Delta^{d})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$. Let us make this assumption as a contradiction hypothesis. Define the function $\mathcal{L}(\mathbf{g}_{n})=\sum_{i=1}^{d}\log(g^{*}(\vartheta_{i})/g_{n}(\vartheta_{i}))g^{*}(\vartheta_{i})$. By the contradiction hypothesis, $\mathcal{L}(\mathbf{g}_{n})(\omega)\rightarrow\infty$ as $n\rightarrow\infty$, for every $\omega\in M$. Then $E^{*}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E^{*}$ denote expectation with respect to $\mathds{P}^{*}$. Let $\overline{\mathbf{g}}_{n}=\eta_{n}\mathbf{g}_{n}+(1-\eta_{n})\mathbf{g}_{n+1}$, with the random variable $\eta_{n}$ taking values in $[0,1]$, be such that
@@
$=\mathcal{L}(\mathbf{g}_{n})-\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})+\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}.$
Notice that
$E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i}\right)=\alpha_{n+1}E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\right),$
@@
$E^{*}\left(\sum_{i=1}^{d}\frac{\mathbf{g}^{*}_{i}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\right)=0,$

### v2

@@ C Proofs for Section 2
Representing probability mass functions on $\Theta_{\eta}$ as vectors in $\Delta_{\eta},$ we can see $\ell(\mathbf{f})=KL(p_{G^{*}}||p_{f})$ as a function on $\Delta_{\eta}$. The assumption (C.6) implies that
$\int_{0}^{\infty}e^{\theta(d_{\eta}^{2}-1)}G^{*}(d\theta)$ $=\int_{0}^{\infty}\sum_{y\in\mathbb{N}_{0}}e^{\log(d_{\eta}^{2})y}e^{-\theta}\frac{\theta^{y}}{y!}G^{*}(d\theta)$ (C.8)
$=\sum_{y\in\mathbb{N}_{0}}e^{2y\log d_{\eta}}p_{G^{*}}(y)<+\infty.$
Thus, for every $i=1,\dots,d_{\eta}$,
$\frac{\partial}{\partial\mathbf{g}_{i}}\ell(\mathbf{g})=\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)\leq e^{\vartheta_{1}-\vartheta_{d_{\eta}}}\sum_{y\in\mathbb{N}_{0}}\left(\frac{\vartheta_{d_{\eta}}}{\vartheta_{1}}\right)^{y}\int_{\Theta}\frac{e^{-\theta}\theta^{y}}{y!}{G^{*}}(\mathrm{d}\theta)\leq e^{\vartheta_{1}-\vartheta_{d_{\eta}}}\int e^{\theta(d_{\eta}-1)}G^{*}(\mathrm{d}\theta)<+\infty$
by (C.8) Analogously, for every $i=1,\dots,d_{\eta}$,
$\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})^{2}}{p_{\mathbf{g}}(y)^{2}}p_{G^{*}}(y)\leq e^{\vartheta_{1}-\vartheta_{d_{\eta}}}\int e^{\theta(d_{\eta}^{2}-1)}G^{*}(\mathrm{d}\theta)<+\infty.$ (C.9)
Thus, $\ell(\mathbf{g})$ is twice continuously differentiable on $\Delta^{d}$. Denoting by $k_{y}(\vartheta_{i})=k(y\mid\vartheta_{i})$ for $i=1,\dots,\vartheta_{d_{\eta}}$, we can write
$\nabla_{\mathbf{g}}KL(p_{G^{*}}||p_{\mathbf{g}})=-\sum_{y\in\mathbb{N}_{0}}\frac{k_{y}-k(y\mid\vartheta_{d_{\eta}})}{p_{\mathbf{g}}(y)}p_{G^{*}}(y).$
The $(i,j)$ element of the Hessian matrix $\mathbf{M}$ of $\ell(\mathbf{g})$ is
$\mathbf{M}_{i,j}$ $=\sum_{y\in\mathbb{N}_{0}}\frac{(k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d_{\eta}}))(k(y\mid\vartheta_{j})-k(y\mid\vartheta_{d_{\eta}}))}{p_{\mathbf{g}}(y)^{2}}p_{G^{*}}(y).$
for $i,j=1,\dots,d-1$.
To prove (C.7), we apply the method of Lagrange multipliers for the constrained minimum of twice continuously differentiable functions. The solution $\mathbf{g}_{\eta}^{*}$ satisfies, for same $\lambda_{i}\geq 0$ ($i=1,\dots,d_{\eta}$) with $\lambda_{i}>0$ for at least one $i$,
$\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d_{\eta}})}{p_{\mathbf{g}_{\eta}^{*}}(y)}p_{G^{*}}(y)-\lambda_{i}+\lambda_{d_{\eta}}=0\quad(i=1,\dots,d_{\eta}-1),$ (C.10)
and
$\lambda_{1}g_{\eta}^{*}(\vartheta_{1})=0,\;\dots,\;\lambda_{d_{\eta}-1}g_{\eta}^{*}(\vartheta_{d_{\eta}-1})=0,\;\lambda_{d_{\eta}}(\sum_{i=1}^{d_{\eta}-1}g_{\eta}^{*}(\vartheta_{i})-1)=0.$ (C.11)
Multiplying (C.10) by $g_{\eta}^{*}(\vartheta_{i})$ and taking the sum with respect to $i=1,\dots,d_{\eta}-1$, we obtain by (C.11)
$\lambda_{d_{\eta}}=-\sum_{i=1}^{d_{\eta}-1}\sum_{y\in\mathbb{N}_{0}}g_{\eta}^{*}(\vartheta_{i})\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)+(1-g_{\eta}^{*}(\vartheta_{d_{\eta}}))\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{d_{\eta}})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y).$
Thus
$\lambda_{d_{\eta}}-\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{d_{\eta}})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)=-\sum_{i=1}^{d_{\eta}}\sum_{y\in\mathbb{N}_{0}}g_{\eta}^{*}(\vartheta_{i})\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)=-1.$
Substituting in (C.10), we obtain
$\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{\mathbf{g}_{\eta}^{*}}(y)}p_{G^{*}}(y)-\lambda_{i}-1=0\quad(i=1,\dots,d_{\eta}-1),$
Thus, by (C.11),
$g_{\eta}^{*}(\vartheta_{i})\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)-1\right)=0$ (C.12)
for every $i=1,\dots,d_{\eta}-1$. Since
$\sum_{i=1}^{d_{\eta}}\sum_{y\in\mathbb{N}_{0}}g_{\eta}^{*}(\vartheta_{i})\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g_{\eta}^{*}}(y)}p_{G^{*}}(y)=1,$
then (C.12) holds also for $i=d_{\eta}$.
$\square$
@@
Under the assumptions of Theorem 2.4, as $n\rightarrow\infty$ the stochastic approximation in $\Delta_{\eta}=\{\mathbf{v}=(\mathbf{v}_{1},\dots,\mathbf{v}_{d_{\eta}-1}):\mathbf{v}_{i}\geq 0,\sum_{i=1}^{d_{\eta}-1}\mathbf{v}_{i}\leq 1\}$:
@@
with
$\mathbf{h}(\mathbf{g})=\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)-\bf 1\right),$
${\bm{\epsilon}}_{n+1}=\mathbf{g}_{n}\circ\left(\frac{\mathbf{k}_{Y_{n+1}}}{p_{\mathbf{g}_{n}}(Y_{n+1})}-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}_{n}}(y)}p_{G^{*}}(y)\right),$
and initial state $\mathbf{g}_{0}\in\mathring{\Delta}^{d_{\eta}}$, converges to $\mathbf{g}_{\eta}^{*}$, $\mathds{P}_{G^{*}}$-a.s.
@@
Let $J$ be the set of vectors in $\Delta^{d_{\eta}}$ that satisfy (C.7). According to Theorem 2.2.1 in Chen (2005), if the following conditions hold:
  • S1
@@
  • S2
    $J$ is closed and there exists a is continuously differentiable function $\mathscr{l}:\Delta^{d_{\eta}}\rightarrow\mathbb{R}$ such that $\mathscr{l}(J)$ is nowhere dense in $\mathbb{R}$ and $\sup_{\delta_{1}\leq{\rm d}(\mathbf{g},J)\leq\delta_{2}}\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}<0$ for every $\delta_{2}>\delta_{1}>0$;
  • S3
    $\lim_{T\rightarrow\infty}\limsup_{n\rightarrow\infty}\frac{1}{T}||\sum_{i=n}^{m(n,T)}\alpha_{i+1}\epsilon_{i+1}||=0,$ $\mathds{P}_{G^{*}}$-a.s., where $m(n,T)=\max\{m:\sum_{i=n}^{m}\alpha_{i+1}\leq T\}$;
  • S4
    $\mathbf{h}$ is measurable and locally bounded,
then
$d(\mathbf{g}_{n},J)\rightarrow 0,$ (C.14)
$\mathds{P}_{G^{*}}$-a.s., as $n\rightarrow\infty$, where $d$ denotes the Euclidean distance in $\mathbb{R}^{d_{\eta}}$.
S1 holds by assumption; S3 holds since $({\bm{\epsilon}}_{n})$ is a martingale difference, and H4 is obvious.
We now prove S2. Let
$\mathscr{l}(\mathbf{g})=\sum_{y\in\mathbb{N}_{0}}\log\left(\frac{p_{G^{*}}(y)}{p_{\mathbf{g}}(y)}\right)p_{G^{*}}(y).$
By (C.6), the function $\mathscr{l}(\mathbf{g})=KL(p_{G^{*}}||p_{\mathbf{g}})$ is twice continuously differentiable and strictly convex in $\Delta^{d_{\eta}}$, as proved in Lemma C.3, with gradient
$\nabla_{\mathbf{g}}\mathscr{l}(\mathbf{g})=-\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}-\mathbf{k}_{y}(\vartheta_{d_{\eta}})\mathbf{1}}{p_{g}(y)}p_{G^{*}}(y).$
The set $J$ is closed. We now prove that $\mathscr{l}(J)$ is nowhere dense in $\mathbb{R}$. The set $J$ is made of a finite number of isolated points. Indeed, let $\mathbf{g}$ be a point in $J$ and let $N=\{i\in\{1,\dots,{d_{\eta}}\}:\mathbf{g}_{i}=0\}$. We prove that the conditions in (C.7) identify the values of $\mathbf{g}_{i}$ for $i\not\in N$. The property is obvious if $N$ contains ${d_{\eta}}-1$ points. Otherwise, assume by contradiction that there exist $\mathbf{g}^{\prime}$ and $\mathbf{g}^{\prime\prime}$ with $\mathbf{g}^{\prime}\neq\mathbf{g}^{\prime\prime}$ such that $\mathbf{g}_{i}^{\prime}=\mathbf{g}_{i}^{\prime\prime}=0$ for $i\in N$ and $\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}}(y)}=1$ for every $i\not\in N$ both for $\mathbf{g}=\mathbf{g}^{\prime}$ and for $\mathbf{g}=\mathbf{g}^{\prime\prime}$. Then, for every $a\in(0,1)$
$1=\sum_{i\not\in N}(a\mathbf{g}^{\prime}_{i})+(1-a)\mathbf{g}^{\prime\prime}_{i})\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{ap_{\mathbf{g}^{\prime}}(y)+(1-a)p_{\mathbf{g}^{\prime\prime}}(y)}$
$<\sum_{i\not\in N}(a\mathbf{g}^{\prime}_{i}+(1-a)\mathbf{g}^{\prime\prime}_{i})\left(a\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}^{\prime}}(y)}+(1-a)\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}^{\prime\prime}}(y)}\right)$
@@
which leads to a contradiction. Thus $J$ is made of a finite number of isolated points. It follows that $\{\mathscr{l}(\mathbf{g}):\mathbf{g}\in J\}$ is nowhere dense in $\mathbb{R}$. To conclude the proof of S2, note that, for $\mathbf{g}\in J^{c}$
$\mathbf{h}(\mathbf{g})^{T}\nabla_{\mathbf{g}}\mathscr{l}$ $=-\left(\mathbf{g}\circ\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)-\bf 1\right)\right)^{T}\left(\sum_{y\in\mathbb{N}_{0}}\frac{\mathbf{k}_{y}-\mathbf{k}_{y}(\vartheta_{d_{\eta}})\mathbf{1}}{p_{\mathbf{g}}(y)}p_{G^{*}}(y)\right)$
$=-\left(\sum_{i=1}^{d_{\eta}}\mathbf{g}_{i}\left(\sum_{y\in\mathbb{N}_{0}}k_{y}(\vartheta_{i})\frac{p_{G^{*}}(y)}{p_{\mathbf{g}}(y)}-1\right)^{2}\right)<0.$
Hence S2 holds, and, by (C.14)
$\inf_{\mathbf{g}\in J}||\mathbf{g}_{n}-\mathbf{g}||\rightarrow 0,$
$\mathds{P}_{G^{*}}$-a.s. as $n\rightarrow\infty$.
We now prove that $\mathbf{g}_{n}$ converges $\mathds{P}$-a.s. to the minimizer $\mathbf{g}_{\eta}^{*}$ of the Kullback-Leibler divergence $KL(p_{G^{*}}||p_{g})$.
First, notice that, since $\mathbf{g}_{n+1}-\mathbf{g}_{n}$ converges to zero $\mathds{P}$-a.s. and since $J$ is made of isolated points, then $\mathbf{g}_{n}$ converges $\mathds{P}$-a.s. to a random vector taking values in $J$.
The minimizer $\mathbf{g}_{\eta}^{*}$ of $\mathscr{l}(\mathbf{g})$ satisfies
$(\mathbf{g}_{\eta}^{*}-\mathbf{g})^{T}\nabla\mathscr{l}(\mathbf{g})<0$
for every $\mathbf{g}\neq\mathbf{g}_{\eta}^{*}$. Hence
$\sum_{i=1}^{{d_{\eta}}-1}(\mathbf{g}_{\eta,i}^{*}-\mathbf{g}_{i})\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})-k(y\mid\vartheta_{d_{\eta}})}{p_{g}(y)}p_{G^{*}}(y)>0,$
which implies that
$\sum_{i=1}^{{d_{\eta}}}\mathbf{g}_{i}^{*}\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g}(y)}p_{G^{*}}(y)-1\right)>0,$
or, equivalently
$\sum_{i=1}^{{d_{\eta}}}(\mathbf{g}_{i}^{*}-\mathbf{g}_{i})\left(\sum_{y\in\mathbb{N}_{0}}\frac{k(y\mid\vartheta_{i})}{p_{g}(y)}p_{G^{*}}(y)-1\right)>0.$ (C.15)
We now show that that, for every $\delta>0$, $\mathds{P}_{G^{*}}(\liminf_{n}{\rm d}(\mathbf{g}_{n},J_{\delta})=0)=0$, where $J_{\delta}=\{\mathbf{g}\in J:d(\mathbf{g},\mathbf{g}_{\eta}^{*})>\delta\}$.
If $\mathbf{g}\in J_{\delta}$, then there exists $\vartheta_{i}$ such that $g(\vartheta_{i})=0$ and $g_{\eta}^{*}(\vartheta_{i})\neq 0$. By contradiction, suppose that there exists $M$ with $\mathds{P}_{G^{*}}(M)>0$ such that ${\rm d}(\mathbf{g}_{n}(\omega),J_{\delta})\rightarrow 0$ for every $\omega\in M$, as $n\rightarrow\infty$. Define the function $\mathcal{L}(\mathbf{g}_{n})=\sum_{i=1}^{d_{\eta}}\log(g_{\eta}^{*}(\vartheta_{i})/g_{n}(\vartheta_{i}))g_{\eta}^{*}(\vartheta_{i})$. By the contradiction hypothesis, $\mathcal{L}(\mathbf{g}_{n})(\omega)\rightarrow\infty$ as $n\rightarrow\infty$, for every $\omega\in M$. Then $E_{G^{*}}({\mathcal{L}}(\mathbf{g}_{n}))\rightarrow\infty$ as $n\rightarrow+\infty$, where $E_{G^{*}}$ denote expectation with respect to $\mathds{P}_{G^{*}}$. Let $\overline{\mathbf{g}}_{n}=\eta_{n}\mathbf{g}_{n}+(1-\eta_{n})\mathbf{g}_{n+1}$, with the random variable $\eta_{n}$ taking values in $[0,1]$, be such that
@@
$=\mathcal{L}(\mathbf{g}_{n})-\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})+\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{(\overline{\mathbf{g}}_{n,i})^{2}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})^{2}.$
Denoting by $\mathcal{G}_{n}$ the sigma-algebra generated by $Y_{1},\dots,Y_{n}$, we can write that
$E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{g}_{n+1,i}-\mathbf{g}_{n,i})\mid\mathcal{G}_{n}\right)=\alpha_{n+1}E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}(\mathbf{h}_{i}(\mathbf{g}_{n})+{\bm{\epsilon}}_{n+1,i})\mid\mathcal{G}_{n}\right),$
@@
$E_{G^{*}}\left(\sum_{i=1}^{d_{\eta}}\frac{\mathbf{g}_{\eta,i}^{*}}{\mathbf{g}_{n,i}}{\bm{\epsilon}}_{n+1,i}\mid\mathcal{G}_{n}\right)=0,$